I came out tonight–the first clear night in almost two months–to observe some variable stars and take several images of standard star fields.
Unfortunately, the filter wheel wouldn’t work (ERROR=1) even after replacing the O-ring and cleaning the new O-ring.
The technical manual suggests there can be a problem with grounding if some of the RJ45 (ethernet) cable pins into the controller are bent, but they do not seem to be.
Next step will be to call Optec support and see what they say, and/or open up the filter housing and clean the motor shaft.
But, for now at least, the telescope is not usable.